5.0 mole of ammonia were introduced into a 5.0 L reaction chamber in which it is partially decomposed at high temperature. At equilibrium at a particular temperature,80% of the ammonia had reacted. calculate Kc?

1
Expert's answer
2022-02-20T08:36:14-0500

The decomposition reaction is below:

2NH3 --> N2 + 3H2

The initial concentration of NH3 = "\frac{5.0\ mol}{5.0\ L}=1.0\ M"

Knowing that 80% of NH3 had reacted, the final concentration of NH3 is 0.20 M. Based on this, an ICE table can be constructed:



Therefore,


"K_c=\frac{[N_2][H_2]^3}{[NH_3]^2}=\frac{0.40\times(1.2)^3}{(0.20)^2}\approx17"


Answer: 17